# Extending Quantum-Safe Communications to Real-World Networks: An Adaptive Security Framework

### v1

@@ Abstract
The advent of quantum computing threats classical cryptographic mechanisms, demanding new strategies for securing communication networks. Since real-world networks cannot be fully Quantum Key Distribution (QKD)-enabled due to infrastructure constraints, practical security solutions must support hybrid operation. This paper presents an adaptive security framework that enables quantum-safe communications across real-world heterogeneous networks by combining QKD and Post-Quantum Cryptography (PQC). Building upon a hierarchical key management architecture with Virtual Key Management Systems (vKMS) and a centralized Quantum Security Controller (QuSeC), the framework dynamically assigns security levels based on node capabilities. By transitioning between pure QKD, hybrid, and PQC modes, it ensures end-to-end quantum-safe protection regardless of the underlying node capabilities. The framework has been implemented and validated on a Kubernetes-based containerized testbed, demonstrating robust operation and performance across all scenarios. Results highlight its potential to support the gradual integration of quantum-safe technologies into existing infrastructures, paving the way toward fully quantum-safe communication networks.
@@ I INTRODUCTION
The development of quantum computing introduces a paradigm shift in the field of cybersecurity. While quantum computers promise revolutionary advancements in several fields, they also pose a significant threat to classical cryptographic algorithms. In particular, widely used public-key algorithms such as RSA and ECC are considered vulnerable to quantum attacks, potentially compromising the asymmetric primitives that underpin key exchange and digital signatures in modern communication systems [easttom2022quantum]. This renders current mechanisms for establishing and managing cryptographic keys insecure.
To address these threats, two primary approaches have been developed for achieving quantum-safe key establishment. The first approach is Post-Quantum Cryptography (PQC), which replaces classical public-key algorithms with alternatives based on hard mathematical problems believed to resist quantum attacks. PQC algorithms can be implemented using existing classical hardware infrastructure, making them easily deployable across current networks. However, their security ultimately depends on computational assumptions [jenefa2023pqc].
The second approach is Quantum Key Distribution (QKD), which avoids public-key cryptography entirely by generating symmetric keys between two legitimate parties using the principles of quantum mechanics. Unlike PQC, QKD provides Information-Theoretic Security (ITS), guaranteed by physical laws rather than computational hardness assumptions [nadlinger2022experimental].
However, several challenges hinder its widespread adoption within current communication infrastructures. First, QKD requires specialized quantum hardware, which increases deployment cost and complexity [diamanti2016practical]. Second, the achievable distance is limited by the lack of quantum repeaters, restricting QKD to metropolitan-scale links. Third, extending QKD beyond point-to-point connections often requires trusted relay nodes, which may weaken the security guarantees by introducing additional points of vulnerability. Fourth, many QKD systems require dedicated dark fibers for quantum signal transmission, complicating integration with existing optical infrastructures and limiting scalability.
Although a fully QKD-enabled network represents the ideal quantum-safe communication paradigm, current constraints make such deployments impractical today. Instead, real-world networks are usually heterogeneous, comprising both QKD-enabled and classical nodes that rely on PQC, as depicted in Fig. 1. Each node is defined as a site that may include one or more pieces of network equipment, quantum equipment, or other related infrastructure. Hybrid architectures that strategically integrate both approaches offer a realistic path forward: they enable the gradual incorporation of QKD while ensuring continuity of service through PQC. This heterogeneity, however, introduces new challenges for providing end-to-end (E2E) quantum-safe communication across mixed networks.
@@
Building on our previous work on scalable QKD architectures [sanz2025quantumsafescalablenetworksopen], which introduced a hierarchical Key Management System (KMS) with Virtual KMSs (vKMS) and a centralized Quantum Security Controller (QuSeC) for efficient KMS discovery and multi-hop relay path computation, this paper extends the model to heterogeneous networks. While that earlier solution improved scalability with compliance with ETSI specifications, it assumed universal QKD capability, limiting its applicability in transitional or hybrid environments.
@@
  2. 2.
    Unified Architecture: An adaptive security framework that extends quantum security to heterogeneous real-world networks by combining QKD and PQC under centralized management, ensuring tailored security for all nodes. The framework also applies communication network concepts to operate in a efficient, scalable and resilient form. The framework leverages communication network principles to operate efficiently, scalably, and resiliently. It also extends the vKMS + QuSeC architecture to orchestrate security policies, assignments, and interoperability across QKD and PQC domains.
@@
The rest of the paper is organized as follows: Section II reviews prior work on QKD and PQC integration, and on extending quantum-safe security to heterogeneous networks. Then, Section III summarizes our previous work on E2E quantum key establishment in QKD networks, on which this proposal builds. Subsequently, Section IV introduces the adaptive security framework, describing the system architecture and its operation. Section V presents the implementation, testbed setup, and validation results. Finally, Section VI presents the main conclusions of the work.
@@ II RELATED WORK
Some theoretical foundations of hybrid Key Encapsulation Mechanisms (KEMs) are formalized in [bindel2019hybrid], where the security of several combiners, such as XOR-then-MAC (xTM) [giacon2018kem], dual-PRF [bellare2024symmetric], and the N Nested Dual-PRF, is introduced and proved. Their results can be used as the foundation for later works that integrate multiple key sources into secure session keys.
@@
Building on this work, [ricci2024hybrid] proposes a three-key combiner that takes as input 256-bit keys from different sources: a classical key generated via Diffie-Hellman, a PQC key generated with KYBER, and a QKD key. They implement and validate the proposal in an FPGA platform, and they also prove IND-CCA security. Similarly, [aquina2024quantum] introduces a quantum secure communication system leveraging a PRF-then-XOR split-key combiner, showing that security can be guaranteed even if only one of the underlying primitives remains secure.
@@
Several works have demonstrated the feasibility of such combiners in realistic cryptosystems. For instance, [zeng2024practical] defines a protocol for symmetric key establishment in QKD–PQC networks between two nodes, considering that several key distribution channels are available between them. The final key is derived through a hybrid KDF applied to multiple distribution channels. The obtained results validate the hybrid protocol where QKD and PQC interoperate within a joint quantum-classical network. In addition, [garms2024experimental] presents Muckle++, a modular cryptosystem integrating hybrid authentication and hybrid key exchange. The proposed protocol derives the secure keys from combining QKD, PQC, and currently used public-key cryptography using a HMAC-based KDF (HKDF). The protocol is validated through real QKD hardware, FPGA-based PQC-KEM acceleration on an FPGA, and server-side implementations. Along similar lines, [doring2022post] integrates PQC key exchange mechanisms into a QKD platform, enabling interoperability between the two technologies and providing a means to deliver quantum-safe keys to nodes where QKD is not available. They demonstrate the feasibility of the proposed scheme in the Berlin OpenQKD testbed, enabling quantum-safe communication between network nodes and extending the solution to wireless, last-mile, and end-user applications. These contributions highlight that hybrid protocols are both practical and deployable, bridging the gap between theoretical combiners and operational systems.
@@
The role of KMSs in hybrid architectures has also been explored in detail. For instance, [james2023key] provides a comprehensive analysis of KMS requirements for large-scale QKD networks, classical and quantum KMSs and emphasizing the need for E2E guarantees even in the presence of compromised trusted nodes. In addition, to strengthen the overall security guarantees provided by a QKD network, they leverage PQC-based relaying of QKD keys between end nodes, with the final session key derived via hybridization at the KMS level. In a complementary direction, [brauer2024linking] demonstrates several approaches for achieving quantum-safe key exchange between three QKD testbeds in Europe, Berlin, Madrid, and Poznan, leveraging PQC-based encryption to emulate long-range QKD links. Different key exchange methods were defined to enable E2E communication between all the nodes of all the participating networks. In addition, they define a key hybridization process by applying a KDF as an internal process on the KMS that manages different internal key stores. These works illustrate how KMSs can serve as the central point for orchestrating hybridization in large-scale deployments.
@@
These efforts primarily focus on extending or strengthening fully QKD-enabled networks and providing a solid foundation for the combination of several input keys through KDFs. However, some works also address environments where only a subset of nodes are QKD-capable. In this context, [zhu2023qkd] introduces the concept of partially-QKD-deployed optical networks (pQKD-ON), where only part of the optical nodes are deployed with QKD systems, and define the problem of key provisioning for secure E2E services in such networks. This work highlights the core challenge of enabling secure E2E communication in heterogeneous settings. They propose a multi-level key pool slicing scheme where QKD key pools are sliced and some portions of keys are pushed outward via lightweight secure transfer. This approach enables having pre-shared QKD key-pairs between non-QKD-deployed nodes. Although they address key provisioning in heterogeneous networks, their approach lacks the centralized management and dynamic security level adaptation that our framework provides. Their solution focuses on static key pool slicing without the unified control plane and or standards compliance that enables real-world deployment.
@@
III BACKGROUND: E2E QUANTUM KEY ESTABLISHMENT IN QKD NETWORKS
@@
This hierarchical design supports both direct key delivery within direct QKD endpoints and multi-hop trusted relay across multiple nodes. The framework ensures compliance with ETSI specifications, including ETSI GS QKD 014 [ETSI_GS_QKD_014] for key delivery, ETSI GS QKD 015 [ETSI_GS_QKD_015] for SDN integration, and the ETSI GS QKD 020 [ETSI_GS_QKD_020], which is under development, for KMS interoperability in relay procedures.
@@
IV PROPOSED FRAMEWORK
This section presents an adaptive framework that extends quantum-safe protection to all nodes within a network. The framework employs dynamic security assignment, adapting the level of security to the specific characteristics and capabilities of each node. In this context, the centralized network and security management plays a crucial role, providing a comprehensive view of the network to enable efficient and robust security management.
@@ IV-A System Architecture
The proposed architecture comprises two different node types. On the one hand, QKD-enabled Nodes (QN) are defined as nodes with access to QKD infrastructure capable of generating and exchanging QKD keys with peer QNs. Each QN hosts at least one KMS responsible for secure storage, synchronization, and delivery of QKD keys on demand. In Fig. 2, nodes D, E, and F are categorized as QN.
On the other hand, Classical Nodes (CN) lack QKD infrastructure, but are capable of employing PQC algorithms to achieve secure communication. In the example architecture shown in Fig. 2, nodes A, B, and C represent the CN category.
Considering this, the proposed architecture is structured into several communication planes, as depicted in Fig. 2:
@@
  • Quantum Key Management Plane: This plane comprises the KMSs responsible for managing QKD keys. Each QKD module in the Quantum Plane is paired with a KMS. The main functions of the KMS include secure storage and management of QKD keys, on demand delivery of QKD keys to authorized applications, and synchronization with the peer KMS.
@@ IV-B Operational description of the security framework
The adaptive security framework ensures quantum-safe E2E communication by dynamically assigning Security Levels based on the involved nodes. The framework is designed to always provide the highest Security Level achievable given the network capabilities and node types involved in the communication.
@@
The complete E2E process operates through six sequential steps that transform an application’s key request into a secure, quantum-safe session key. Fig. 3 depicts the high-level operational flow of the proposed framework, illustrating how these steps interact across the different architectural planes. The process begins when an application initiates a secure session request and concludes with the delivery of quantum-safe keys to both communicating endpoints.
@@
    Secure Session Request. An initiating application requests a quantum-safe communication session with another application to its associated vKMS. This request is made sending a Get Key message defined in the ETSI GS QKD 014 API. The vKMS receives this request and prepares to coordinate with the QuSeC to deliver the secure key.
@@
    Secure Key Derivation. This is the execution phase where all configured components work together to generate the shared session key. The specific operations depend on the assigned Security Level: QKD-enabled nodes may retrieve quantum keys from their KMSs, relay operations may forward keys through intermediate trusted nodes using OTP encryption, PQC operations may establish shared secrets using post-quantum KEMs, and hybrid approaches may combine multiple key sources through cryptographic KDFs. Throughout this process, each vKMS performs its assigned role while maintaining synchronization with its peer to ensure both endpoints derive the identical session key.
@@
After both applications have received the secure keys, they can proceed with communication using the established session key. The framework is designed to be protocol-agnostic, allowing the quantum-safe key to be integrated into different security mechanisms, such as pre-shared keys, session keys, or encryption keys, depending on the requirements of the underlying communication protocol.
@@
For the assignment of a Security Level, each vKMS, upon receiving a secure key request from an application, queries the QuSeC to request a Security Level for the communication between the initiating and the target applications. The proposed framework adapts the Security Level based on whether the involved applications belong to QN or CN, and on the quantum path availability. The QuSeC assigns a Security Level to each E2E communication and coordinates all the associated vKMSs and KMSs to execute the required operations to derive a symmetric E2E key with the highest security possible. In this context, after analyzing all possible scenarios, four different Security Levels have been defined:
  • Security Level 1 - Direct QN-to-QN. This Security Level applies when two applications, each belonging to QNs directly connected via a QKD link, request secure communication. The assigned Security Level is ITS, as no classical key exchange is involved, thereby fully preserving the quantum-safe properties of QKD. Fig. 4(a) shows the high-level operation of Level 1, between applications on nodes E and F.
  • Security Level 2 - Multi-hop QN-to-QN. This Security Level is assigned when two applications on QNs without a direct QKD link require secure communication. In such cases, the QKD with a trusted relay approach is employed, where keys are forwarded through intermediate trusted nodes leveraging One-Time Pad (OTP) encryption. Within the declared trust boundary for relays and with secure control and API channels, secrecy arguments are information-theoretic for the key material. Fig. 4(b) illustrates this Security Level for applications on nodes D and F.
  • Security Level 3 - CN-to-QN. This level applies when one application resides on a node without QKD infrastructure (CN). In this case, a relay and hybridization approach is employed to ensure secure key generation. Specifically, a hardened KDF derives a session key from at least one QKD-derived segment and one PQC KEM shared secret. The resulting key remains secure as long as at least one input key is secure. This multi-source approach produces a quantum-resistant symmetric key by combining independent security mechanisms. Fig. 4(c) depicts the high-level operation of Level 3 between applications on nodes C and D, including the relay of a QKD key via node E and the hybridization process.
  • Security Level 4 - CN-to-CN. This Security Level applies when both communicating applications reside on CNs, meaning no QKD infrastructure is available. In this case, a purely PQC-based approach is employed to ensure secure key exchange and derivation. Additionally, a KDF hybridization approach can be leveraged by combining multiple cryptographic sources to further enhance resilience. While this level does not provide ITS as in QKD-based approaches, it ensures robust protection against both classical and quantum threats. Fig. 4(d) shows this Security Level for applications on nodes A and B.
@@
After level assignment, the vKMS sends a level-based request to the QuSeC, which computes a level-specific configuration that encodes the session context, the selected cryptographic suites, and the minimal set of instructions required by each participant. For Levels 1 and 2, this step has been described in detail in our previous work, and it primarily involves KMS identification and, when applicable, relay path installation. For Levels 3 and 4, however, additional structure is required.
@@
Although each participant has a distinct role, all of them receive the same session configuration message. This message consists of common session information, such as a unique identifier, along with a policy block specific to each participant. The policy block determines the actions to be performed, the selected cryptographic suites, and the identities of the endpoints involved.
@@
For Level 3, the role-scoped steps that must be performed are described as follows, and they are graphically represented in Fig. 5:

### v2

# Ane Sanz    Eire Salegi    Asier Atutxa    David Franco    Jasone Astorga    Eduardo Jacob

@@ Abstract
The advent of quantum computing threatens classical cryptographic mechanisms, demanding new strategies for securing communication networks. Since real-world networks cannot be fully Quantum Key Distribution (QKD)-enabled due to infrastructure constraints, practical security solutions must support hybrid operation. This paper presents an adaptive security framework that enables quantum-safe communications across real-world heterogeneous networks by combining QKD and Post-Quantum Cryptography (PQC). Building upon a hierarchical key management architecture with Virtual Key Management Systems (vKMS) and a centralized Quantum Security Controller (QuSeC), the framework dynamically assigns security levels based on node capabilities. By transitioning between pure QKD, hybrid, and PQC modes, it ensures end-to-end quantum-safe protection regardless of the underlying node capabilities. The framework has been implemented and validated on a Kubernetes-based containerized testbed, demonstrating robust operation and performance across all scenarios. Results highlight its potential to support the gradual integration of quantum-safe technologies into existing infrastructures, paving the way toward fully quantum-safe communication networks.
@@ I Introduction
The development of quantum computing introduces a paradigm shift in the field of cybersecurity, with direct implications for quantum communication and quantum cryptography. While quantum computers promise revolutionary advancements in several fields, they also pose a significant threat to classical cryptographic algorithms. In particular, widely used public-key algorithms such as RSA and ECC are considered vulnerable to quantum attacks, potentially compromising the asymmetric primitives that underpin key exchange and digital signatures in modern communication systems [9]. This renders current mechanisms for establishing and managing cryptographic keys insecure.
To address these threats, two primary approaches have been developed for achieving quantum-safe key establishment. The first approach is Post-Quantum Cryptography (PQC), which replaces classical public-key algorithms with alternatives based on hard mathematical problems believed to resist quantum attacks. PQC algorithms can be implemented using existing classical hardware infrastructure, making them easily deployable across current networks. However, their security ultimately depends on computational assumptions [13].
The second approach is Quantum Key Distribution (QKD), which avoids public-key cryptography entirely by generating symmetric keys between two legitimate parties using the principles of quantum mechanics. Unlike PQC, QKD provides Information-Theoretic Security (ITS), guaranteed by physical laws rather than computational hardness assumptions [14].
However, several challenges hinder its widespread adoption within current communication infrastructures. First, QKD requires specialized quantum hardware, which increases deployment cost and complexity [6]. Second, the achievable distance is limited by the lack of quantum repeaters, restricting QKD to metropolitan-scale links. Third, extending QKD beyond point-to-point connections often requires trusted relay nodes, which may weaken the security guarantees by introducing additional points of vulnerability. Fourth, many QKD systems require dedicated dark fibers for quantum signal transmission, complicating integration with existing optical infrastructures and limiting scalability.
Although a fully QKD-enabled network represents the ideal quantum-safe communication paradigm, current constraints make such deployments impractical today. Instead, real-world networks are usually heterogeneous, comprising both QKD-enabled and classical nodes that rely on PQC, as depicted in Figure 1. Each node is defined as a site that may include one or more pieces of network equipment, quantum equipment, or other related infrastructure. Hybrid architectures that strategically integrate both approaches offer a realistic path forward: they enable the gradual incorporation of QKD while ensuring continuity of service through PQC. This heterogeneity, however, introduces new challenges for providing end-to-end (E2E) quantum-safe communication across mixed networks.
@@
Our previous work on scalable QKD architectures [19] introduced a hierarchical Key Management System (KMS) with Virtual KMSs (vKMS) and a centralized Quantum Security Controller (QuSeC) for efficient KMS discovery and multi-hop relay path computation. Building on that work, this paper extends the model to heterogeneous networks. While that earlier solution improved scalability with compliance with ETSI specifications, it assumed universal QKD capability, limiting its applicability in transitional or hybrid environments.
@@
    Unified Architecture: An adaptive security framework that extends quantum security to heterogeneous real-world networks by combining QKD and PQC under centralized management, ensuring tailored security for all nodes. The framework also applies communication network concepts to operate in an efficient, scalable, and resilient form. The framework leverages communication network principles to operate efficiently, scalably, and resiliently. It also extends the vKMS + QuSeC architecture to orchestrate security policies, assignments, and interoperability across QKD and PQC domains.
  2. 2.
@@
The rest of the paper is organized as follows: Section II reviews prior work on QKD and PQC integration, and on extending quantum-safe security to heterogeneous networks. Then, Section III summarizes our previous work on E2E quantum key establishment in QKD networks, on which this proposal builds. Subsequently, Section IV introduces QuLore, the adaptive security framework, describing the system architecture and its operation. Section V summarizes the security analysis of the framework, and Section VI presents the implementation, testbed setup, and performance results. Finally, Section VII presents the main conclusions of the work.
@@ II Related work
Some theoretical foundations of hybrid Key Encapsulation Mechanisms (KEMs) are formalized in [4], where the security of several combiners, such as XOR-then-MAC (xTM) [11], dual-PRF [3], and the N Nested Dual-PRF, is introduced and proved. Their results can be used as the foundation for later works that integrate multiple key sources into secure session keys.
@@
Building on this work, [18] proposes a three-key combiner that takes as input 256-bit keys from different sources: a classical key generated via Diffie-Hellman, a PQC key generated with KYBER, and a QKD key. They implement and validate the proposal in an FPGA platform, and they also prove IND-CCA security. Similarly, [2] introduces a quantum secure communication system leveraging a PRF-then-XOR split-key combiner, showing that security can be guaranteed even if only one of the underlying primitives remains secure.
@@
Several works have demonstrated the feasibility of such combiners in realistic cryptosystems. For instance, [20] defines a protocol for symmetric key establishment in QKD–PQC networks between two nodes, considering that several key distribution channels are available between them. The final key is derived through a hybrid KDF applied to multiple distribution channels. The obtained results validate the hybrid protocol where QKD and PQC interoperate within a joint quantum-classical network. In addition, [10] presents Muckle++, a modular cryptosystem integrating hybrid authentication and hybrid key exchange. The proposed protocol derives the secure keys from combining QKD, PQC, and currently used public-key cryptography using a HMAC-based KDF (HKDF). The protocol is validated through real QKD hardware, FPGA-based PQC-KEM acceleration, and server-side implementations. Along similar lines, [8] integrates PQC key exchange mechanisms into a QKD platform, enabling interoperability between the two technologies and providing a means to deliver quantum-safe keys to nodes where QKD is not available. They demonstrate the feasibility of the proposed scheme in the Berlin OpenQKD testbed, enabling quantum-safe communication between network nodes and extending the solution to wireless, last-mile, and end-user applications. These contributions highlight that hybrid protocols are both practical and deployable, bridging the gap between theoretical combiners and operational systems.
@@
The role of KMSs in hybrid architectures has also been explored in detail. For instance, our previous work [19] proposed and validated a key management framework for QKD networks leveraging virtual KMSs under centralized control. Since that work provides the foundation for the approach presented here, it is explained in detail in Section III. Additionally, [12] provides a comprehensive analysis of KMS requirements for large-scale QKD networks, classical and quantum KMSs, and emphasizes the need for E2E guarantees even in the presence of compromised trusted nodes. In addition, to strengthen the overall security guarantees provided by a QKD network, they leverage PQC-based relaying of QKD keys between end nodes, with the final session key derived via hybridization at the KMS level. In a complementary direction, [5] demonstrates several approaches for achieving quantum-safe key exchange between three QKD testbeds in Europe, Berlin, Madrid, and Poznan, leveraging PQC-based encryption to emulate long-range QKD links. Different key exchange methods were defined to enable E2E communication between all the nodes of all the participating networks. In addition, they define a key hybridization process by applying a KDF as an internal process on the KMS that manages different internal key stores. These works illustrate how KMSs can serve as the central point for orchestrating hybridization in large-scale deployments. These efforts primarily focus on extending or strengthening fully QKD-enabled networks and providing a solid foundation for the combination of several input keys through KDFs.
@@
Additionally, some works also address environments where only a subset of nodes are QKD-capable. In this context, [21] introduces the concept of partially-QKD-deployed optical networks (pQKD-ON), where only part of the optical nodes are deployed with QKD systems, and defines the problem of key provisioning for secure E2E services in such networks. This work highlights the core challenge of enabling secure E2E communication in heterogeneous settings. They propose a multi-level key pool slicing scheme where QKD key pools are sliced, and some portions of keys are pushed outward via lightweight secure transfer. This approach enables having pre-shared QKD key-pairs between non-QKD-deployed nodes. Although they address key provisioning in heterogeneous networks, their approach lacks the centralized management and dynamic security level adaptation that our framework provides. Their solution focuses on static key pool slicing without the unified control plane and or standards compliance that enables real-world deployment.
@@
III Background: E2E quantum-safe key establishment in QKD networks
@@
This hierarchical design supports both direct key delivery within direct QKD endpoints and multi-hop trusted relay across multiple nodes. The framework ensures compliance with ETSI specifications, including ETSI GS QKD 014 [17] for key delivery, ETSI GS QKD 015 [15] for SDN integration, and the ETSI GS QKD 020 [16], which is under development, for KMS interoperability in relay procedures.
@@
IV QuLore: a framework for quantum-safe key establishment
This section presents QuLore [^1] QuLore is inspired by eguzkilore, a traditional Basque flower that is hung on house doors as a protective symbol. In Basque mythology, the eguzkilore represents the light of the sun, guarding the house and keeping witches and evil spirits away from home.: an adaptive framework that extends quantum-safe protection to all nodes within a network. The framework employs dynamic security assignment, adapting the level of security to the specific characteristics and capabilities of each node. In this context, the centralized network and security management plays a crucial role, providing a comprehensive view of the network to enable efficient and robust security management.
@@ IV-A System Architecture
The proposed architecture comprises two different node types. On the one hand, QKD-enabled Nodes (QN) are defined as nodes with access to QKD infrastructure capable of generating and exchanging QKD keys with peer QNs. Each QN hosts at least one KMS responsible for secure storage, synchronization, and delivery of QKD keys on demand. In Figure 2, nodes D, E, and F are categorized as QN.
On the other hand, Classical Nodes (CN) lack QKD infrastructure, but are capable of employing PQC algorithms to achieve secure communication. In the example architecture shown in Figure 2, nodes A, B, and C represent the CN category.
Considering this, the proposed architecture is structured into several communication planes, as depicted in Figure 2:
@@
  • Quantum Key Management Plane: This plane comprises the KMSs responsible for managing QKD keys. Each QKD module in the Quantum Plane is paired with a KMS. The main functions of the KMS include secure storage and management of QKD keys, on-demand delivery of QKD keys to authorized applications, and synchronization with the peer KMS.
@@ IV-B Operational description of the security framework
The adaptive security framework ensures quantum-safe E2E communication by dynamically assigning Security Levels based on the involved nodes. The QuLore framework is designed to always provide the highest Security Level achievable, given the network capabilities and node types involved in the communication.
@@
The complete E2E process operates through six sequential steps that transform an application’s key request into a secure, quantum-safe session key. Figure 3 depicts the high-level operational flow of the proposed framework, illustrating how these steps interact across the different architectural planes. The process begins when an application initiates a secure session request and concludes with the delivery of quantum-safe keys to both communicating endpoints.
@@
    Secure Session Request. An initiating application requests a quantum-safe communication session with another application to its associated vKMS. This request is made by sending a Get Key message defined in the ETSI GS QKD 014 API. The vKMS receives this request and prepares to coordinate with the QuSeC to deliver the secure key.
@@
    Secure Key Derivation. This is the execution phase where all configured components work together to generate the shared session key. The specific operations depend on the assigned Security Level: purely QKD-based operations may involve retrieving quantum keys from their KMSs, relay operations may forward keys through intermediate trusted nodes using OTP encryption, PQC operations may establish shared secrets using post-quantum KEMs, and hybrid approaches may combine multiple key sources through cryptographic KDFs. Throughout this process, each vKMS performs its assigned role while maintaining synchronization with its peer to ensure both endpoints derive the identical session key.
@@
After both applications have received the secure keys, they can proceed with communication using the established session key. The QuLore framework is designed to be protocol-agnostic, allowing the quantum-safe key to be integrated into different security mechanisms, such as pre-shared keys, session keys, or encryption keys, depending on the requirements of the underlying communication protocol.
@@
Each vKMS, upon receiving a secure key request from an application, queries the QuSeC to request a Security Level for the communication between the initiating and the target applications. The proposed framework adapts the Security Level based on whether the involved applications belong to QN or CN, and on the quantum path availability. The QuSeC assigns a Security Level to each E2E communication and coordinates all the associated vKMSs and KMSs to execute the required operations to derive a symmetric E2E key with the highest security possible. In this context, after analyzing all possible scenarios, four different Security Levels have been defined:
  • Security Level 1 - Direct QN-to-QN. This Security Level applies when two applications, each belonging to QNs directly connected via a QKD link, request secure communication. The assigned Security Level is ITS, as no classical key exchange is involved, thereby fully preserving the quantum-safe properties of QKD. Figure 4(a) shows the high-level operation of Level 1, between applications on nodes E and F.
  • Security Level 2 - Multi-hop QN-to-QN. This Security Level is assigned when two applications on QNs without a direct QKD link require secure communication. In such cases, the QKD with a trusted relay approach is employed, where keys are forwarded through intermediate trusted nodes leveraging One-Time Pad (OTP) encryption. Within the declared trust boundary for relays and with secure control and API channels, secrecy arguments are information-theoretic for the key material. Figure 4(b) illustrates this Security Level for applications on nodes D and F.
  • Security Level 3 - CN-to-QN. This level applies when one application resides on a node without QKD infrastructure (CN). In this case, a relay and hybridization approach is employed to ensure secure key generation. Specifically, a hardened KDF derives a session key from at least one QKD-derived segment and one PQC KEM shared secret. The resulting key remains secure as long as at least one input key is secure. This multi-source approach produces a quantum-resistant symmetric key by combining independent security mechanisms. Figure 4(c) depicts the high-level operation of Level 3 between applications on nodes C and D, including the relay of a QKD key via node E and the hybridization process.
  • Security Level 4 - CN-to-CN. This Security Level applies when both communicating applications reside on CNs, meaning no QKD infrastructure is available. In this case, a purely PQC-based approach is employed to ensure secure key exchange and derivation. Additionally, a KDF hybridization approach can be leveraged by combining multiple cryptographic sources to further enhance resilience. While this level does not provide ITS as in QKD-based approaches, it ensures robust protection against both classical and quantum threats. Figure 4(d) shows this Security Level for applications on nodes A and B.
@@
After level assignment, the vKMS sends a level-based request to the QuSeC, which computes a level-specific configuration that encodes the session context, the selected cryptographic suites, and the minimal set of instructions required by each participant. For Levels 1 and 2, this step has been described in detail in our previous work [19], and it primarily involves KMS identification and, when applicable, relay path installation. For Levels 3 and 4, however, additional structure is required.
@@
Although all participants receive a session configuration message with the same overall structure, the content of the message is not identical for every participant. Each message includes common session information, such as a unique session identifier, as well as a policy block tailored to the role of the recipient. This policy block specifies the actions to be performed, the selected cryptographic suites, and the identities of the relevant endpoints.
@@
For Level 3, the role-scoped steps that must be performed are described as follows, and they are graphically represented in Figure 5: